OpenStudy (anonymous):

TRE OR FALSE. The sum of three consecutive integers is 36. The three integers are 11, 12, and 13.

OpenStudy (anonymous):

11 + 12 +13 = 36

OpenStudy (anonymous):

x + x+1 + x+2 = 36 3x + 3 = 36 x + 1 = 12 x = 12 -1 = 11 -> x +1 = 12 -> x+ 2 = 13